# Unsupervised Attributed Dynamic Network Embedding with Stability Guarantees

### v1

@@ Abstract
Stability for dynamic network embeddings ensures that nodes behaving the same at different times receive the same embedding, allowing comparison of nodes in the network across time. We present attributed unfolded adjacency spectral embedding (AUASE), a stable unsupervised representation learning framework for dynamic networks in which nodes are attributed with time-varying covariate information. To establish stability, we prove uniform convergence to an associated latent position model. We quantify the benefits of our dynamic embedding by comparing with state-of-the-art network representation learning methods on three real attributed networks. To the best of our knowledge, AUASE is the only attributed dynamic embedding that satisfies stability guarantees without the need for ground truth labels, which we demonstrate provides significant improvements for link prediction and node classification.
@@ 1 Introduction
Approaches that purposefully exploit node attributes include frameworks based on matrix factorisation [Liu et al., 2020, Li et al., 2017], deep learning [Tang et al., 2022, Ahmed et al., 2024, Wei et al., 2019], or Bayesian modelling [Luodi et al., 2024]. However, to the best of our knowledge, none of these methods have stability guarantees, which ensure that if two node/time pairs “behave the same” in the network, their representation is the same up to noise. Stability allows for the comparison of embeddings over time because the embedding space has a consistent interpretation.
@@
Figure 1 shows the embeddings of AUASE and two comparable methods visually explaining the difference between stable and unstable embeddings. The astute reader will be wondering why leading dynamic graph methods could have such a visually poor internal representation. There are a wide class of problems - for which graph neural networks (GNNs) were developed - that can be expressed as a supervised learning problem, and the classifier learns a separate map for each time point, for instance, node classification and link prediction [Rossi et al., 2020, Pareja et al., 2020]. However, many problems are not of this form and require unsupervised stable embeddings. Examples include but are not limited to: group discovery in purchasing networks for marketing purposes [Palla et al., 2007], anomaly detection in financial or cybersecurity networks [Ranshous et al., 2015] and protein function prediction in dynamic biological networks [Klein et al., 2012, Yue et al., 2020].
We compare AUASE to state-of-the-art unsupervised attributed dynamic embeddings on the tasks of node classification and link prediction as a means to demonstrate AUASE embedding’s validity and the value of stability. In extensive experiments on three large real-world datasets AUASE considerably outperforms other methods, which in many cases perform worse than baseline guessing precisely because they lack embedding stability.
@@ 2.2 Attributed Dynamic Network Model
Conditional on latent positions $Z_{i}^{(t)}\in\mathcal{Z}$ and fixed latent positions $Z_{n+\ell}^{(t)}=i_{\ell}$, the sequence of symmetric matrices $\mathbf{A}_{C}^{(1)},\ldots,\mathbf{A}_{C}^{(T)}\in\mathbb{R}^{(n+p)\times(n+p)}$ is distributed as an attributed dynamic latent position network model with sparsity parameter $\rho>0$, if, for all $i<j$,
@@ 2.3 Theoretical Results
There exists maps $\phi:\mathcal{Z}^{T}\cup\mathcal{I}^{T}\rightarrow\mathbb{R}^{d}$ and $\phi_{t}:\mathcal{Z}\cup\mathcal{I}\rightarrow\mathbb{R}^{d}$ for all $t\in[T]$ such that
@@
$\mathbf{P}\left(\lvert f_{l}(Z_{i}^{(t)})-\mathbb{E}[f_{l}(Z_{i}^{(t)})]\rvert>x\mid Z_{i}^{(t)}\right)\leq\textnormal{exp}\left(-\beta x\right)$
@@
Assumption 2 is a technical condition which ensures that the growth of the singular values of $\mathbf{P}_{C}$ is regulated. Assumption 3 assumes a fixed number of covariates which alongside Assumption 4 ensures that the networks are dense enough to recover the latent positions. Versions of these assumptions are required to analyse the multilayer random dot product graph [Jones and Rubin-Delanchy, 2021]. These assumptions are not limiting as for finite $n$ (i.e. any practical application) scaling $\alpha$ would mitigate any imbalance between sparsity in the graph and the covariates.
@@ 2.4 Stability Properties
We say that node/time pairs $(i,s)$ and $(j,t)$ are exchangeable if they “behave the same” within the dynamic network. For the attributed dynamic latent position model (Definition 3) this is equivalent to stating, for all $Z\in\mathcal{Z}\cup\mathcal{I}$,
@@ 3.2 Implementation and Efficiency
Both AUASE and UASE can be easily implemented with the Spectral-Embedding Python package [^1] https://github.com/iggallagher/Spectral-Embedding. The computational times to compute the dynamic embeddings of three large real-data networks for each baseline are reported in Table 2. The runtime of AUASE implemented with the Spectral-Embedding Python package is orders of magnitude faster than competing methods, showing that AUASE is highly efficient for large graphs. For larger networks, AUASE can be implemented highly efficiently using the irlba package in R or irlbpy in Python, which quickly computes leading singular vectors [^2] https://bwlewis.github.io/irlba/comparison.html.
@@ 3.3 Simulated Example
where the latent position index into the community edge probability matrix $\mathbf{B}\in\mathbb{R}^{3\times 3}$ and mean attribute matrix $\mathbf{D}\in\mathbb{R}^{3\times p}$,
@@
We construct the attributed dynamic latent position model with $n=1000$ nodes and $p=150$ covariates over $T=10$ time points. Nodes are assigned with equal probability to three communities, depicted by the colours in Figure 2a, corresponding to a trajectory of local behaviours $(Z^{(1)}_{i},\dots,Z^{(T)}_{i})$, the shapes in Figure 2a. The edge probabilities are $p_{0}=0.5$, $p_{1}=0.3$ and mean attributes $\bm{\mu}_{1},\bm{\mu}_{2}\in\mathbb{R}^{p}$ are detailed in Appendix C.
@@
Figure 2c-2f shows the dynamic network embedding for these four techniques. For visualisation, we reduce the three-dimensional embeddings to one dimension with UMAP [McInnes et al., 2020]. The solid line shows the mean embedding for each community with a 90% confidence interval.
@@ 3.4 Datasets
As discussed in Section 2.4, stability is a desirable intrinsic property of a dynamic embedding essential for any downstream analysis. To provide an objective measure of its value, we evaluate the learned node representations on the downstream tasks node classification and link prediction for the following three attributed dynamic networks. Further details about the construction of the adjacency matrices $\mathbf{A}^{(t)}$ and covariate matrices $\mathbf{C}^{(t)}$ are given in Appendix D.
  • DBLP [^3] https://www.aminer.org/citation [Ley, 2009]: Co-authorship network consisting of bibliography data from computer science publications. Nodes represent authors with edges denoting co-authorship. Attributes are derived from an author’s abstracts published in each time interval. For each interval an author has one of seven labels based on their most frequent publication area.
@@
Table 1: Attributed dynamic network statistics for the DBLP, ACM and Epinions datasets showing the number of nodes $n$, the number of covariates $p$, the number of time intervals $T$, and the number of node labels $K$.
@@
For the methods listed in Section 3.1, the embedding dimension is selected using ScreeNOT [Zhu and Ghodsi, 2006] giving $d=15$, $d=29$ and $d=22$ for DBLP, ACM and Epinions, respectively. The one exception is DySAT on the ACM dataset where $d=32$. We use degree correction for AUASE and UASE embeddings [Passino et al., 2022].
@@ 3.5 Node Classification
Node classification aims to predict label categories for current nodes using the historical node embedding data. We train a classifier using XGBoost [Chen and Guestrin, 2016] on the first 65% of time intervals for DBLP, ACM and Epinions and test on the remaining data. We reserve 10% of the training data for validation to select the weight hyperparameters via cross-validation for AUASE, DRLAN and CONN. Other hyperparameters are set to the default values provided by the original authors.
Figure 3: Node label classification accuracies for DBLP, ACM and Epinions datasets for each time interval in the corresponding test datasets. The dynamic embedding techniques outlined in Section 3.1 are solid lines, while a basic baseline predicting the most common label is dashed.
Figure 3 reports the classification accuracy for each test year, while Table 2 reports the mean accuracy over all test data. Other metrics were considered and gave equivalent results (see Appendix D, Tables 4, 5 and 6). Because label prediction requires predicting current embeddings, the stability illustrated in Figure 1 allows AUASE to dramatically outperform baselines in terms of accuracy - 40% over DyRep on DBLP, 23% over DySAT on ACM, and 11% over GloDyNE on Epinions.
AUASE also performs noticeably better than UASE suggesting that the inclusion of covariate information is improving the quality of the dynamic embedding. Stability clearly is driving this performance increase, as Figure 3 shows that no unstable dynamic method consistently outperforms the baseline classifier of predicting the most common class.
@@
Method Task DBLP ACM Epinions
@@
Node Class. 0.219 0.526 0.131
Time $>$1h $>$1h $>$1h
@@
Node Class. 0.185 0.568 0.175
Time $\approx$10m $>$1h $\approx$1h
DRLAN Link Pred. $0.575\pm 0.018$ $0.641\pm 0.015$ $0.626\pm 0.081$
Node Class. 0.232 0.573 0.118
Time $<$1s $<$1s $<$1s
@@
Node Class. 0.272 0.578 0.143
Time $>$5h $>$40h $>$48h
@@
Node Class. 0.175 0.591 0.095
Time $>$1h $>$5h $>$3h
UASE Link Pred. $0.907$±$\pm 0.003$ $\underline{0.896}\pm 0.003$ $\underline{0.806}\pm 0.009$
Node Class. 0.55 0.758 0.201
Time $<$1s $<$5s $<$1s
AUASE Link Pred. $\mathbf{0.915}\pm 0.004$ $\underline{0.896}\pm 0.005$ $\mathbf{0.809}\pm 0.009$
Node Class. 0.668 0.825 0.281
Time $<$5s $<$5s $<$5s
@@ 3.6 Link Prediction
These are assigned a positive label if there exists an edge between the two nodes at time $t+1$, namely, $\mathbf{A}_{ij}^{(t+1)}=1$, otherwise, assigned a negative label, balanced so the classes have the same size. We train a binary classifier using XGBoost [Chen and Guestrin, 2016] on the first $T-2$ intervals for DBLP, ACM and Epinions, test on the remaining data and repeat the sampling 10 times.
@@ 3.7 Parameter Sensitivity
In an unsupervised setting, there are many proposed heuristics for hyperparameter estimation, e.g. Tsitsulin et al. [2023], which might result in different choices of $\alpha$ in AUASE. Importantly, the information contained in the embedding is not sensitive to this choice. To demonstrate this, Figure 4 presents the prediction accuracy for node classification on AUASE embeddings of the DBLP dataset.
@@ 4 Conclusion
By combining provable stability over time with attributed embedding, we have shown that AUASE learns intrinsically ‘better’ dynamic attributed embeddings that solve problems that all previous (i.e. unstable) methods cannot. This empirically improves predictive performance for node classification and link prediction problems, it’s expected to help with unsupervised tasks and aids interpretability.
@@ C.2 Hyperparameter selection
We set the embedding dimension $d=3$ for all the methods, the known number of communities. The selection of hyperparameters for each method shown in Figure 2 is outlined below:
@@
  • DRLAN - visually chosen parameters: $\beta=0.7$, $p=2$, $q=1$, $\alpha=[1,0.1,0.001,0.0001]$, $\theta=[1,1,0.1,0.001,0.0001]$.

### v2

@@ Abstract
Stability for dynamic network embeddings ensures that nodes behaving the same at different times receive the same embedding, allowing comparison of nodes in the network across time. We present attributed unfolded adjacency spectral embedding (AUASE), a stable unsupervised representation learning framework for dynamic networks in which nodes are attributed with time-varying covariate information. To establish stability, we prove uniform convergence to an associated latent position model. We quantify the benefits of our dynamic embedding by comparing with state-of-the-art network representation learning methods on four real attributed networks. To the best of our knowledge, AUASE is the only attributed dynamic embedding that satisfies stability guarantees without the need for ground truth labels, which we demonstrate provides significant improvements for link prediction and node classification.
@@ 1 Introduction
Approaches that purposefully exploit node attributes include frameworks based on matrix factorisation [Liu et al., 2020, Li et al., 2017], deep learning [Tang et al., 2022, Ahmed et al., 2024, Wei et al., 2019], or Bayesian modelling [Luodi et al., 2024]. However, to the best of our knowledge, none of these methods have stability guarantees, which ensure that if two node/time pairs ‘behave the same’ in the network, their representation is the same up to noise. Stability allows for the comparison of embeddings over time because the embedding space has a consistent interpretation.
@@
Figure 1 shows the embeddings of AUASE and two comparable methods - visually explaining the difference between stable and unstable embeddings. The astute reader will be wondering why leading dynamic graph methods could have such a visually poor internal representation. There are a wide class of problems - for which graph neural networks (GNNs) were developed - that can be expressed as a supervised learning problem, and the classifier learns a separate map for each time point, for instance, node classification and link prediction [Rossi et al., 2020, Pareja et al., 2020]. However, many problems are not of this form and require unsupervised stable embeddings. Examples include but are not limited to: group discovery in purchasing networks for marketing purposes [Palla et al., 2007], anomaly detection in financial or cybersecurity networks [Ranshous et al., 2015] and protein function prediction in dynamic biological networks [Klein et al., 2012, Yue et al., 2020].
We compare AUASE to state-of-the-art unsupervised attributed dynamic embeddings on the tasks of node classification and link prediction as a means to demonstrate AUASE embedding’s validity and the value of stability. In extensive experiments on four large real-world datasets AUASE considerably outperforms other methods, which in many cases perform worse than baseline guessing precisely because they lack embedding stability.
@@ 2.2 Attributed Dynamic Network Model
Conditional on latent positions $Z_{i}^{(t)}\in\mathcal{Z}$ and fixed latent positions $Z_{n+\ell}^{(t)}=i_{\ell}$, the sequence of symmetric matrices $\mathbf{A}_{C}^{(1)},\ldots,\mathbf{A}_{C}^{(T)}\in\mathbb{R}^{(n+p)\times(n+p)}$ is distributed as an attributed dynamic latent position network model with sparsity parameter $\rho>0$, if,
for all $i<j$,
@@ 2.3 Theoretical Results
There exist maps $\phi:\mathcal{Z}^{T}\cup\mathcal{I}^{T}\rightarrow\mathbb{R}^{d}$ and $\phi_{t}:\mathcal{Z}\cup\mathcal{I}\rightarrow\mathbb{R}^{d}$ for all $t\in[T]$ such that
@@
$\mathbb{P}\left(\lvert f_{l}(Z_{i}^{(t)})-\mathbb{E}[f_{l}(Z_{i}^{(t)})]\rvert>x\mid Z_{i}^{(t)}\right)\leq\textnormal{exp}\left(-\beta x\right)$
@@
Assumption 2 is a technical condition which ensures that the growth of the singular values of $\mathbf{P}_{C}$ is regulated. Assumption 3 assumes a fixed number of covariates which alongside Assumption 4 ensures that the networks are dense enough to recover the latent positions. Versions of these assumptions are required to analyse the multilayer random dot product graph [Jones and Rubin-Delanchy, 2021]. These assumptions are not limiting as for finite $n$ (i.e., any practical application) scaling $\alpha$ would mitigate any imbalance between sparsity in the graph and the covariates.
@@ 2.4 Stability Properties
We say that node/time pairs $(i,s)$ and $(j,t)$ are exchangeable if they ‘behave the same’ within the dynamic network. For the attributed dynamic latent position model (Definition 3) this is equivalent to stating, for all $Z\in\mathcal{Z}\cup\mathcal{I}$,
@@ 3.2 Implementation and Efficiency
Both AUASE and UASE can be easily implemented with the pyemb Python package [^1] https://pyemb.github.io/pyemb/html/index.html. The code for the reproducing experiments is available at this GitHub repository  [^2] https://github.com/emmaceccherini/AUASE.
The computational times to compute the dynamic embeddings of four large real-data networks for each baseline are reported in Table 2. The runtime of AUASE implemented with the pyemb Python package is orders of magnitude faster than competing methods, showing that AUASE is highly efficient for large graphs.
We do not provide a comprehensive time and space complexity analysis as AUASE is essentially a sparse truncated SVD whose space and time complexity has been extensively studied. The space complexity in the sparsest regime we consider is $O(Tn\log^{k}(n))$.
The truncated SVD of $\mathbf{A}_{C}$ can be computed using the Augmented Implicitly Restarted Lanczos Bidiagonalization algorithm [Baglama and Reichel, 2005] implemented in the irlba package in R, or the irlbpy in Python. Although the exact time complexity of this algorithm has not been studied theoretically, it is known that the time complexity of more general Lanczos-type algorithms for computing the rank truncated SVD is $O(Nd)$ [Tomás et al., 2023], in the sparsest regime we consider, this is $O(Tn\log^{k}(n)d)$. The author of the irlba package demonstrates its speed by performing a simulated experiment in which they compute the first 2 singular vectors of a sparse 10M x 10M matrix with 1M non-zero entries which takes approximately 6 seconds on a computer with two Intel Xeon CPU E5-2650 processors (16 physical CPU cores) running at 2 GHz equipped with 128 GB of ECC DDR3 RAM  [^3] https://bwlewis.github.io/irlba/comparison.html.
@@ 3.3 Simulated Example
where the latent position indexes into the community edge probability matrix $\mathbf{B}\in\mathbb{R}^{3\times 3}$ and the mean attribute matrix $\mathbf{D}\in\mathbb{R}^{3\times p}$,
@@
We construct the attributed dynamic latent position model with $n=1000$ nodes and $p=150$ covariates over $T=10$ time points. Nodes are assigned with equal probability to three communities, depicted by the colours in Figure 2a, corresponding to a trajectory of local behaviours $(Z^{(1)}_{i},\dots,Z^{(T)}_{i})$, represented by the shapes in Figure 2a. The edge probabilities are $p_{0}=0.5$, $p_{1}=0.3$ and mean attributes $\bm{\mu}_{1},\bm{\mu}_{2}\in\mathbb{R}^{p}$ are detailed in Appendix C.
@@
Figures 2c-2f show the dynamic network embedding for these four techniques. For visualisation, we reduce the three-dimensional embeddings to one dimension with UMAP McInnes et al. [2020]. The solid line shows the mean embedding for each community with a 90% confidence interval.
@@ 3.4 Datasets
As discussed in Section 2.4, stability is a desirable intrinsic property of a dynamic embedding essential for any downstream analysis. To provide an objective measure of its value, we evaluate the learned node representations on the downstream tasks node classification and link prediction for the following four attributed dynamic networks. Further details about the construction of the adjacency matrices $\mathbf{A}^{(t)}$ and covariate matrices $\mathbf{C}^{(t)}$ are given in Appendix D.
  • DBLP [^4] https://www.aminer.cn/citation [Ley, 2009]: Co-authorship network consisting of bibliography data from computer science publications. Nodes represent authors with edges denoting co-authorship. Attributes are derived from an author’s abstracts published in each time interval. For each interval an author has one of seven labels based on their most frequent publication area.
@@
  • ogbn-mag  [^6] https://ogb.stanford.edu/docs/nodeprop [Wang et al., 2020]: Co-authorship network from the Microsoft Academic Graph. We construct 10 adjacency matrices and attributes matrices where authors are the nodes. The tasks we perform are at author-level not paper-level (usual for ogbn-mag) which creates much harder tasks (see Appendix D for more details).
@@
Table 1: Attributed dynamic network statistics for the datasets showing the number of nodes $n$, the number of covariates $p$, the number of time intervals $T$, and the number of node labels $K$.
@@
ogbn-mag 479979 128 10 349
The embedding dimension is selected using ScreeNOT [Zhu and Ghodsi, 2006] giving $d=15$, $d=29$ and $d=22$ for DBLP, ACM and Epinions, respectively. The one exception is DySAT on the ACM dataset where $d=32$. For ogbn-mag we choose $d=50$ based on the scree plot. We use degree correction for AUASE and UASE embeddings [Passino et al., 2022].
The number of nodes in ogbn-mag is an order of magnitude larger than in the other datasets, therefore only UASE, AUASE and DRLAN are computationally feasible. More details about the computational barriers of DySAT, GloDyNE, DyRep and CONN are given in Appendix D.
@@
Figure 3: Node label classification accuracies for DBLP, ACM, Epinions and ogbn-mag datasets for each time interval in the corresponding test datasets. The dynamic embedding techniques outlined in Section 3.1 are solid lines, and different colors represents different methods. A basic baseline predicting the most common label is dashed.
@@ 3.5 Node Classification
Node classification aims to predict label categories for current nodes using the historical node embedding data. We train a classifier using XGBoost [Chen and Guestrin, 2016] on the first 65% of time intervals for all the datasets and test on the remaining data. We reserve 10% of the training data for validation to select the weight hyperparameters via cross-validation for AUASE, DRLAN and CONN. Other hyperparameters are set to the default values provided by the original authors.
Figure 3 reports the classification accuracy for each test year, while Table 2 reports the mean accuracy over all test data. Other metrics were considered and gave equivalent results (see Appendix D, Tables 4, 5 and 6). Because label prediction requires predicting current embeddings, the stability illustrated in Figure 1 allows AUASE to dramatically outperform baselines in terms of accuracy - 40% over DyRep on DBLP, 23% over DySAT on ACM, 11% over GloDyNE on Epinions and 12% over DRLAN on ogbn-mag. AUASE also performs noticeably better than UASE, suggesting that the inclusion of covariate information is improving the quality of the dynamic embedding. Stability is clearly driving this performance increase, as Figure 3 shows that no unstable dynamic method consistently outperforms the baseline classifier of predicting the most common class.
@@
Method Task DBLP ACM Epinions ogbn-mag
@@
Node Class. 0.219 0.526 0.131 -
Time $\approx$1h $\approx$1h $\approx$1h -
@@
Node Class. 0.185 0.568 0.175 -
Time $\approx$10m $\approx$1h $\approx$1h -
DRLAN Link Pred. $0.575\pm 0.018$ $0.641\pm 0.015$ $0.626\pm 0.081$ $\underline{0.939}\pm 0.009$
Node Class. 0.232 0.573 0.118 0.044
Time $\approx$1s $\approx$1s $\approx$1s $\approx$ 10m
@@
Node Class. 0.272 0.578 0.143 -
Time $\approx$5h $\approx$40h $\approx$48h -
@@
Node Class. 0.175 0.591 0.095 -
Time $\approx$1h $\approx$5h $\approx$3h -
UASE Link Pred. $0.907$±$\pm 0.003$ $\underline{0.896}\pm 0.003$ $\underline{0.806}\pm 0.009$ $0.911\pm 0.002$
Node Class. 0.550 0.758 0.201 0.140
Time $\approx$1s $\approx$5s $\approx$1s $\approx$30m
AUASE Link Pred. $\mathbf{0.915}\pm 0.004$ $\underline{0.896}\pm 0.005$ $\mathbf{0.809}\pm 0.009$ $\mathbf{0.954}\pm 0.003$
Node Class. 0.668 0.825 0.281 0.173
Time $\approx$5s $\approx$5s $\approx$5s $\approx$30m
@@ 3.6 Link Prediction
These are assigned a positive label if there exists an edge between the two nodes at time $t+1$, namely, $\mathbf{A}_{ij}^{(t+1)}=1$, otherwise, they are assigned a negative label, balanced so the classes have the same size. We train a binary classifier using XGBoost [Chen and Guestrin, 2016] on the first $T-2$ intervals for all the datasets, test on the remaining data and repeat the sampling 10 times.
@@ 3.7 Parameter Sensitivity
In an unsupervised setting, there are many proposed heuristics for hyperparameter estimation, e.g., Tsitsulin et al. [2023], which might result in different choices of $\alpha$ in AUASE. Importantly, the information contained in the embedding is not sensitive to this choice. To demonstrate this, Figure 4 presents the prediction accuracy for node classification on AUASE embeddings of the DBLP dataset.
@@ 4 Conclusion
By combining provable stability over time with attributed embedding, we have shown that AUASE learns intrinsically ‘better’ dynamic attributed embeddings that solve problems that all previous (i.e., unstable) methods cannot. This empirically improves predictive performance for node classification and link prediction problems, it is expected to help with unsupervised tasks and aids interpretability.
@@
Acknowledgements.
Emma Ceccherini gratefully acknowledges support by the Centre for Doctoral Training in Computational Statistics and Data Science (Compass, EPSRC Grant number EP/S023569/1). This work was carried out using the computational facilities of the Advanced Computing Research Centre, University of Bristol - http://www.bris.ac.uk/acrc/. Andrew Jones gratefully acknowledges support by the NeST EPSRC programme grant EP/X002195/1.
@@ C.2 Hyperparameter selection
We set the embedding dimension $d=3$ for all the methods, the known number of communities. Experiments with other embedding dimension are given in https://anonymous.4open.science/r/AUASE-80E4/README.md. The selection of hyperparameters for each method shown in Figure 2 is outlined below:
@@
  • DRLAN - visually chosen parameters (see https://anonymous.4open.science/r/AUASE-80E4/README.md): $\beta=0.7$, $p=2$, $q=1$, $\alpha=[1,0.1,0.001,0.0001]$, $\theta=[1,1,0.1,0.001,0.0001]$.